# Divergence-Based Similarity Function for Multi-View Contrastive Learning

### v2

@@ Abstract
Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of data. While prior methods incorporate multiple views at the loss or feature level, they primarily capture pairwise relationships and fail to model the joint structure across all views. In this work, we propose a divergence-based similarity function (DSF) that explicitly captures the joint structure by representing each set of augmented views as a distribution and measuring similarity as the divergence between distributions. Extensive experiments demonstrate that DSF consistently improves performance across diverse tasks—including kNN classification, linear evaluation, transfer learning, and distribution shift—while also achieving greater efficiency than other multi-view methods. Furthermore, we establish a connection between DSF and cosine similarity, and demonstrate that, unlike cosine similarity, DSF operates effectively without the need for tuning a temperature hyperparameter.
@@ 1 Introduction
  • Feature-level approaches: Aggregating features (e.g., via averaging (Pototzky, Sultan, and Schmidt-Thieme 2022) or optimal transport (Piran et al. 2024)) before applying the contrastive loss.
@@
To address this, we propose a novel similarity function that models the joint structure among all augmented views of data. We estimate a probability distribution on the unit hypersphere and compute similarity via negative divergence between distributions, see Figure 1. We use the von Mises–Fisher (vMF) distribution, a widely used Gaussian-like distribution on the unit hypersphere (Mardia and Jupp 2009; Fisher, Lewis, and Embleton 1993; Kitagawa and Rowley 2022; Scott, Gallagher, and Mozer 2021; Du et al. 2024). vMF distribution is parameterized by mean direction $\mu$ and concentration $\kappa$, which respectively represent the central tendency and joint structure of features. Notably, in the two-view case, our method recovers cosine similarity in the InfoNCE loss (Oord, Li, and Vinyals 2018), linking pairwise and multi-view contrastive frameworks.
@@ 2.1 Original Contrastive Learning
CL appears in many self-supervised learning methods. SimCLR (Chen et al. 2020a) leverages the InfoNCE loss (Oord, Li, and Vinyals 2018) to pull together representations of two augmented views while pushing apart all other samples. MoCo (He et al. 2020) introduces a momentum encoder and a dynamic dictionary (negative queue) to maintain a large and consistent set of negative samples. DCL (Yeh et al. 2022) decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable and high performance training with smaller batches and without a momentum encoder. MoCo v3 (Chen, Xie, and He 2021) explores CL using a transformer-based model, specifically the Vision Transformer (ViT) (Dosovitskiy et al. 2020).
@@ 2.2 Contrastive Learning with $M\geq 3$ views
Existing $M\geq 3$ multi‐view CL methods have broadly fallen into two categories. The first category shapes the loss by averaging pairwise terms across views. (Shidani et al. 2024) generates $M$ views per a sample and aggregates their pairwise losses either arithmetically or geometrically. The second category is a method that deals with view representations at the feature level before computing the final objective. FastSiam (Pototzky, Sultan, and Schmidt-Thieme 2022) averages the other (M1M-1) view features to construct a target of each view, and M3G (Piran et al. 2024) adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost. In contrast to these approaches, DSF integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
@@ 5.1 Experimental Setup
We conduct experiments on two standard benchmarks: CIFAR-10 (Krizhevsky, Hinton et al. 2009) and ImageNet-100, a 100-class subset of ImageNet (Deng et al. 2009). These datasets allow us to evaluate the proposed method on both small- and large-scale visual recognition tasks.
We investigated transfer learning performance on the Food-101 dataset (Bossard, Guillaumin, and Van Gool 2014), CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), FGVC Aircraft (Maji et al. 2013), the Describable Textures Dataset (DTD) (Cimpoi et al. 2014) and Oxford-IIIT Pets (Parkhi et al. 2012).
@@
We follow the MoCo v2 (Chen et al. 2020b) framework, using ResNet-18 as the encoder for CIFAR-10, with a 128-dimensional projection head. For ImageNet-100, we adopt MoCo v3 (Chen, Xie, and He 2021) with a ViT-S/16 backbone to evaluate scalability across both CNN and transformer architectures.
@@ 5.4 More Resource Comparison
In Figure 4, we examine how the number of augmented views affects representation quality, as increasing views generally improves performance (Tian, Krishnan, and Isola 2020; Shidani et al. 2024; Pototzky, Sultan, and Schmidt-Thieme 2022). In DSF, more views lead to more accurate estimation of the vMF parameters $\mu$ and $\kappa$, due to reduced variance. As shown in Figure 4, DSF consistently outperforms other multi-view methods across different numbers of views, clearly demonstrating the superiority of DSF in multi-view representation learning.
@@ C.2 Second Term Calculation
In Eq. (LABEL:eq:kld), the second term cannot be evaluated in closed form using the modified Bessel function $I_{p}(\cdot)$. Therefore, we adopt an asymptotic approximation as given in (DLMF , Eq. (10.41.3)).

### v4

@@ Abstract
Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of data. While prior methods incorporate multiple views at the loss or feature level, they primarily capture pairwise relationships and fail to model the joint structure across all views. In this work, we propose a divergence-based similarity function (DSF) that explicitly captures the joint structure by representing each set of augmented views as a distribution and measuring similarity as the divergence between distributions. Extensive experiments demonstrate that DSF consistently improves performance across diverse tasks—including kNN classification, linear evaluation, transfer learning, and distribution shift—while also achieving greater efficiency than other multi-view methods. Furthermore, we establish a connection between DSF and cosine similarity, and demonstrate that, unlike cosine similarity, DSF operates effectively without the need for tuning a temperature hyperparameter. Code and pretrained models are available at https://github.com/Jeon789/DSF.
@@ 1 Introduction
  • Feature-level approaches: Aggregating features (e.g., via averaging (Pototzky et al. 2022) or optimal transport (Piran et al. 2024)) before applying the contrastive loss.
@@
To address this, we propose a novel similarity function that models the joint structure among all augmented views of data. We estimate a probability distribution on the unit hypersphere and compute similarity via negative divergence between distributions, see Figure 1. We use the von Mises–Fisher (vMF) distribution, a widely used Gaussian-like distribution on the unit hypersphere (Mardia and Jupp 2009; Fisher et al. 1993; Kitagawa and Rowley 2022; Scott et al. 2021; Du et al. 2024). vMF distribution is parameterized by mean direction $\mu$ and concentration $\kappa$, which respectively represent the central tendency and joint structure of features. Notably, in the two-view case, our method recovers cosine similarity in the InfoNCE loss (Oord et al. 2018), linking pairwise and multi-view contrastive frameworks.
@@ 2.1 Original Contrastive Learning
CL appears in many self-supervised learning methods. SimCLR (Chen et al. 2020a) leverages the InfoNCE loss (Oord et al. 2018) to pull together representations of two augmented views while pushing apart all other samples. MoCo (He et al. 2020) introduces a momentum encoder and a dynamic dictionary (negative queue) to maintain a large and consistent set of negative samples. DCL (Yeh et al. 2022) decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable and high performance training with smaller batches and without a momentum encoder. MoCo v3 (Chen et al. 2021) explores CL using a transformer-based model, specifically the Vision Transformer (ViT) (Dosovitskiy et al. 2020).
@@ 2.2 Contrastive Learning with $M\geq 3$ views
Existing $M\geq 3$ multi‐view CL methods have broadly fallen into two categories. The first category shapes the loss by averaging pairwise terms across views. (Shidani et al. 2024) generates $M$ views per a sample and aggregates their pairwise losses either arithmetically or geometrically. The second category is a method that deals with view representations at the feature level before computing the final objective. FastSiam (Pototzky et al. 2022) averages the other (M1M-1) view features to construct a target of each view, and M3G (Piran et al. 2024) adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost. In contrast to these approaches, DSF integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
@@ 5.1 Experimental Setup
We conduct experiments on two standard benchmarks: CIFAR-10 (Krizhevsky et al. 2009) and ImageNet-100, a 100-class subset of ImageNet (Deng et al. 2009). These datasets allow us to evaluate the proposed method on both small- and large-scale visual recognition tasks.
We investigated transfer learning performance on the Food-101 dataset (Bossard et al. 2014), CIFAR-10 and CIFAR-100 (Krizhevsky et al. 2009), FGVC Aircraft (Maji et al. 2013), the Describable Textures Dataset (DTD) (Cimpoi et al. 2014) and Oxford-IIIT Pets (Parkhi et al. 2012).
@@
We follow the MoCo v2 (Chen et al. 2020b) framework, using ResNet-18 as the encoder for CIFAR-10, with a 128-dimensional projection head. For ImageNet-100, we adopt MoCo v3 (Chen et al. 2021) with a ViT-S/16 backbone to evaluate scalability across both CNN and transformer architectures.
@@ 5.4 More Resource Comparison
In Figure 4, we examine how the number of augmented views affects representation quality, as increasing views generally improves performance (Tian et al. 2020; Shidani et al. 2024; Pototzky et al. 2022). In DSF, more views lead to more accurate estimation of the vMF parameters $\mu$ and $\kappa$, due to reduced variance. As shown in Figure 4, DSF consistently outperforms other multi-view methods across different numbers of views, clearly demonstrating the superiority of DSF in multi-view representation learning.
@@ C.2 Second Term Calculation
In Eq. (LABEL:eq:kld), the second term cannot be evaluated in closed form using the modified Bessel function $I_{p}(\cdot)$. Therefore, we adopt an asymptotic approximation as given in (DLMF 2024, Eq. (10.41.3)).